OpenStudy (anonymous):

Solve x + 3y = 9 3x - 3y = -13 Question 7 options: 1) (-1, 10/3 ) 2) (1, -10/3 ) 3) (-1, 3/10 ) 4) (1, -3/10 )

OpenStudy (saifoo.khan):

1) (-1, 10/3 )

jimthompson5910 (jim_thompson5910):

Solve for x to get x = 9-3y Now work with the 2nd equation and plug in x = 9-3y 3x-3y=-13 3(9-3y)-3y=-13 27-9y-3y=-13 -12y=-13-27 -12y = -40 y = -40/(-12) y = 10/3 Now plug this back into x = 9-3y and solve for x x = 9-3y x = 9-3(10/3) x = 9-10 x = -1 So the two solutions are x = -1 and y = 10/3 giving us the ordered pair (-1,10/3)

OpenStudy (saifoo.khan):

Hip, you copied wrong!

jimthompson5910 (jim_thompson5910):

yeah you lost y somewhere...

OpenStudy (saifoo.khan):

Lol.